# RANSAC Back to SOTA: A Two-stage Consensus Filtering for Real-time 3D Registration

### v1

@@ I Introduction
Point cloud registration (PCR) seeks to estimate a six-degree-of-freedom (6-DOF) pose to align point clouds, which is crucial for 3D reconstruction[3, 4, 5, 6], robotic navigation[7, 8, 9, 10], and aerial photogrammetry[11, 12, 13, 14]. Iterative closest point (ICP)[15] pioneers a local registration pipeline that iteratively searches correspondences and minimizes their distances. Unfortunately, it often converges to erroneous local optima under notable viewpoint changes[16, 17, 18]. To overcome this, researchers derive coarse poses from feature correspondences[19, 20], known as correspondence-based PCR. However, sensor noises, object occlusions, and descriptor limitations inevitably lead to outliers, presenting great challenges for registration[21, 22, 23].
Random sample consensus (RANSAC) [24, 25, 26, 27, 28] is the most popular solution to handle correspondences with outliers. It follows a generation-and-selection strategy that iteratively samples correspondences to generate and verify hypotheses. However, the requisite iterations escalate exponentially with the outlier ratio, which renders them far inferior to existing methods in accuracy and efficiency.
Figure 1: Average registration recall and runtime for ETH’s tree sequence. The runtime denotes the average time for a single registration. Both our method and MAC achieve a 100% recall. Remarkably, ours accomplishes this in just 59ms, leading to a three-orders-of-magnitude improvement over MAC’s 331,587ms.
Thus, we propose a two-stage consensus filtering (TCF) incorporating one-point and two-point RANSAC steps before the three-point RANSAC to ensure accuracy and significantly boost efficiency. It begins with a one-point RANSAC to remove correspondence outliers via length constraints. Afterward, a two-point RANSAC evaluates the angle consistency to get a more reliable consensus. This design significantly reduces the iteration number by lowering the sampling dimension and also decreases the outlier ratio. As a result, the three-point RANSAC requires only a few iterations, enhancing registration performance. Following this, three-point RANSAC computes a coarse pose and removes outliers based on transformed correspondence’s distances. Finally, a scale-adaptive Cauchy iteratively reweighted least squares (IRLS) is applied to calculate the optimal pose. As depicted in Fig. 1, our method demonstrates SOTA performance and achieves a speedup of up to three orders of magnitude.
The main contributions are as follows:
@@ II-A Handcrafted PCR
Loss Optimization. This category of methods progressively optimizes error functions to refine the pose. Iterative closest point (ICP)[15] pioneers a local registration pipeline that iteratively searches correspondences and minimizes Euclidean distances. Following this, several works seek to advance ICP from diverse perspectives, such as incorporating the point-to-plane metric[32], probabilistic models[33], and branch-and-bound (BnB) theories[34]. Despite robust results, poor initial guess tends to yield wrong results. Fast global registration (FGR)[4] formulates a dense objective featuring a scaled Geman-McClure estimator as the penalty function and alleviates the impact of local minima via graduated non-convexity. Yang et al. [35] propose a versatile approach for robust estimation that utilizes modern non-minimal solvers, extending the applicability of Black-Rangarajan duality and graduated non-convexity (GNC) to various spatial perception tasks, including mesh registration, shape alignment, and pose graph optimization.
@@ II-B Learning-based PCR
Descriptor Learning. Descriptor learning mainly aims to create sparse descriptions for local 3D patches using a weight-sharing network or produce dense descriptions for the entire point cloud in a single forward pass. The pioneering work 3DMatch[3] converts local 3D patches into volumetric representations and then derives descriptors via a siamese convolutional network. FCGF[38] adopts sparse tensor representation and employs Minkowski convolutional neural networks as the backbone to learn dense features. Several works successively boost the feature descriptiveness via N-tuple loss[39], folding-based encoder[40], and spherical CNNs[41]. Predator[21] adds an overlap-attention block to aggregate contextual information for robust registration in low-overlap scenarios. While they eliminate manual feature design and bolster robustness, substantial labeling, training, and optimization remain time-consuming, particularly for low-cost robot platforms.
@@ IV-A Experimental Setup
Evaluation Metrics. We evaluate all methods via registration recall ($rr$), rotation error ($e_{r}$), and translation error ($e_{t}$):
$e_{t}$ $=||\boldsymbol{t}_{e}-\boldsymbol{t}_{g}||$ (11)
$e_{r}$ $=\arccos\frac{tr(\mathbf{R}_{e}(\mathbf{R}_{g})^{\rm T})-1}{2}$
$rr$ $=\frac{N_{success}}{N_{all}}$
where $tr(\cdot)$ is the matrix trace. $\boldsymbol{t}$ and $\mathbf{R}$ denote translation and rotation, respectively. Subscript $g$ denotes ground-truth values while $e$ represents estimated ones. $N_{success}$ is the count of successfully registered point cloud pairs while $N_{all}$ is the total number. Registration is successful if $e_{r}\leq 5\degree,e_{t}\leq 0.5m$ on the ETH dataset, and $e_{r}\leq 5\degree,e_{t}\leq 1m$ on the KITTI dataset.
@@ IV-B Results on ETH
Recall and Accuracy. Fig. 7 depicts our registration results on the ETH dataset and Table II outlines the registration results. Courtyard, facade, and office have inlier ratios exceeding 5% while challenging arch and trees only have around 1%. Our method matches MAC in the recall, achieving top performance in all five scenarios (88% for arch and 100% for the others). It delivers the best translation accuracy in the arch (0.04m), courtyard (0.04m), and trees (0.03m), the second-best in the facade (0.03m) and office (0.06m). Our method also excels in rotational accuracy, leading in the courtyard (0.02°) and ranking second in the facade (0.11°). Our rotation error exceeds MAC, SC2PCR, and TEASER by 0.05°, 0.09°, and 0.07°in the arch, office, and trees, respectively. These results showcase that our method achieves comparable or superior registration accuracy and recall to mainstream methods.
@@ IV-C Results on KITTI
Recall and Accuracy. Table III summarizes the registration results on the KITTI dataset. Fig. 8 depicts our registration results. Due to the large offset (about 17m) between point clouds, all methods experience degradation. Sequence 01 is a high-speed scenario characterized by significant distortion and feature deterioration, resulting in a low inlier ratio of 2.4%. Despite this, our method achieves the highest recall (65%). It attains the second-highest translation accuracy across all five sequences: 0.25m (01), 0.34m (02), 0.25m (03), 0.21m (06), and 0.25m (09). Our rotation accuracy slightly lags behind SC2PCR and MAC but stays within a maximum deviation of 0.18°. These results underscore that our method excels in challenging environments like highways, exhibiting higher registration recall.
@@ IV-D Boosting Performance with TCF
As illustrated in Table IV, we feed our consensus set into other methods to evaluate the registration performance. The total runtime in the table comprises ours and the baseline methods. Our approach reduces trees’ correspondences from 8,842 to 60, boosting the inlier ratio from 1.2% to 98%. All pipelines achieve 100% recall and stable registration, with deviations limited to 0.27° and 0.05m. TEASER and PCR-99 enhance the recall by 30% and 50%, respectively. TEASER, SC2PCR, MAC, and PCR-99 are 332ms, 9s, 330s, and 10s faster than their original methods, respectively. The correspondences in sequence 01 drop from 3343 to 121 while the inlier ratio jumps from 2.4% to 48%. TEASER, SC2PCR and MAC demonstrate 12%, 15%, and 10% recall improvement, respectively. TEASER, SC2PCR, MAC, and PCR-99 become faster by approximately 70ms, 2s, 30s, and 12s, respectively. The results prove that cascading our method with other methods effectively boosts registration performance and facilitates a qualitative leap in speed.
@@
IV-E Ablation Study
@@
TABLE V: Ablation experiments on outlier removal for three RANSAC modules. 1R, 2R, and 3R represent one-point, two-point, and three-point RANSAC, respectively. IR is the inlier ratio. IN means the inlier number.
@@ V Conclusion
We propose a two-stage consensus filtering method for correspondence-based registration. Our approach restores the RANSAC family to state-of-the-art (SOTA) performance. By combining length consistency with angle consistency, we effectively eliminate the majority of outliers, thus reducing the requisite iterations by three-point RANSAC. We combine three RASNAC modules and iterative weighted least squares into a complete registration pipeline that accelerates registration without compromising accuracy. Experimental results demonstrate a three-orders-of-magnitude speed enhancement while maintaining high registration quality. Moreover, integrating our method with existing approaches can notably improve efficiency and recall.

### v2

@@ I Introduction
Point cloud registration (PCR) seeks to estimate a six-degree-of-freedom (6-DOF) pose to align point clouds, which is crucial for 3D reconstruction[3, 4, 5, 6], robotic navigation[7, 8, 9, 10], and aerial photogrammetry[11, 12, 13]. Iterative closest point (ICP)[14] pioneers a local registration pipeline that iteratively searches correspondences and minimizes their distances. Unfortunately, it often converges to erroneous local optima under notable viewpoint changes[15, 16]. To overcome this, researchers derive coarse poses from feature correspondences[17, 18], known as correspondence-based PCR. However, sensor noises, object occlusions, and descriptor limitations inevitably lead to outliers, presenting great challenges for registration[19, 20].
Random sample consensus (RANSAC) [21, 22, 23] is the most popular solution to handle correspondences with outliers. It follows a generation-and-selection strategy that iteratively samples correspondences to generate and verify hypotheses. However, the requisite iterations escalate exponentially with the outlier ratio, which renders them far inferior to existing methods in accuracy and efficiency.
Figure 1: Average registration recall and runtime for ETH’s tree sequence. The runtime denotes the average time for a single registration. Both our method and MAC achieve a 100% recall. Remarkably, ours accomplishes this in just 59 ms, leading to a three-orders-of-magnitude improvement over MAC’s 331587 ms.
Thus, we propose a two-stage consensus filtering (TCF) incorporating one-point and two-point RANSAC steps before the three-point RANSAC to ensure accuracy and significantly boost efficiency. It begins with a one-point RANSAC to remove correspondence outliers via length constraints. Afterward, a two-point RANSAC evaluates the angle consistency to get a more reliable consensus. This design significantly reduces the iteration number by lowering the sampling dimension and also decreases the outlier ratio. As a result, the three-point RANSAC requires only a few iterations, enhancing registration performance. Following this, three-point RANSAC computes a coarse pose and removes outliers based on transformed correspondence’s distances. Finally, a scale-adaptive Cauchy iteratively reweighted least squares (IRLS) is applied to calculate the optimal pose. As depicted in Fig. 1, our method demonstrates SOTA performance and achieves a speedup of up to three orders of magnitude. Our contributions are as follows:
@@ II-A Classical PCR
Optimization-based. This category of methods progressively optimizes error functions to refine the pose. Iterative closest point (ICP)[14] pioneers a local registration pipeline that iteratively searches correspondences and minimizes Euclidean distances. Following this, several works seek to advance ICP from diverse perspectives, such as incorporating the point-to-plane metric[27], probabilistic models[28], and branch-and-bound (BnB) theories[29]. Despite robust results, poor initial guess tends to yield wrong results. Fast global registration (FGR)[4] formulates a dense objective featuring a scaled Geman-McClure estimator as the penalty function and alleviates the impact of local minima via graduated non-convexity. Yang et al. [30] propose a versatile approach for robust estimation that utilizes modern non-minimal solvers, extending the applicability of Black-Rangarajan duality and graduated non-convexity to various spatial perception tasks.
@@ II-B Learning-based PCR
Descriptor Learning. Descriptor learning mainly aims to create sparse descriptions for local 3D patches using a weight-sharing network or produce dense descriptions for the entire point cloud in a single forward pass. The pioneering work 3DMatch[3] converts local 3D patches into volumetric representations and then derives descriptors via a siamese convolutional network. FCGF[33] adopts sparse tensor representation and employs Minkowski convolutional neural networks as the backbone to learn dense features. Several works successively boost the feature descriptiveness via N-tuple loss[34], folding-based encoder[35], and spherical CNNs[36]. While they eliminate manual feature design and bolster robustness, substantial labeling, training, and optimization remain time-consuming, particularly for low-cost robot platforms.
@@ IV-A Experimental Setup
Evaluation Metrics. We evaluate all methods via registration recall $rr=\frac{N_{success}}{N_{all}}$, rotation error $e_{r}=\arccos\frac{tr(\mathbf{R}_{e}(\mathbf{R}_{g})^{\rm T})-1}{2}$, and translation error $e_{t}=||\boldsymbol{t}_{e}-\boldsymbol{t}_{g}||$. $tr(\cdot)$ is the matrix trace. $\boldsymbol{t}$ and $\mathbf{R}$ denote translation and rotation, respectively. Subscript $g$ denotes ground-truth values while $e$ represents estimated ones. $N_{success}$ is the count of successfully registered point cloud pairs while $N_{all}$ is the total number. Registration is successful if $e_{r}\leq 5\degree,e_{t}\leq 0.5\ m$ on the ETH dataset, and $e_{r}\leq 5\degree,e_{t}\leq 1\ m$ on the KITTI dataset.
@@ IV-B Results on ETH
Recall and Accuracy. Fig. 7 depicts our registration results on the ETH dataset and Table II outlines the registration results. Courtyard, facade, and office have inlier ratios exceeding 5% while challenging arch and trees only have around 1%. Our method matches MAC in the recall, achieving top performance in all five scenarios. It delivers the best translation accuracy in the arch (0.04 m), courtyard (0.04 m), and trees (0.03 m), the second-best in the facade (0.03 m) and office (0.06 m).It also excels in rotational accuracy, leading in the courtyard (0.02°) and ranking second in the facade (0.11°). Our rotation error exceeds MAC, SC2PCR, and TEASER by 0.05°, 0.09°, and 0.07°in the arch, office, and trees, respectively. These results showcase that our method achieves comparable or superior registration accuracy and recall to mainstream methods.
@@ IV-C Results on KITTI
Recall and Accuracy. Table III summarizes the registration results on KITTI dataset. Fig. 8 depicts our registration results. The large offset (about 17 m) between point clouds degrades performance for all methods. Sequence 01 is a high-speed scenario characterized by significant distortion and feature deterioration, resulting in a low inlier ratio of 2.4%. Despite this, our method achieves the highest recall (65%). It attains the second-highest translation accuracy across all five sequences. Our rotation accuracy slightly lags behind SC2PCR and MAC but stays within a maximum deviation of 0.18°. These results underscore that our method excels in challenging environments like highways, exhibiting higher registration recall.
@@ IV-D Boosting Performance with TCF
As illustrated in Table IV, we feed our consensus set into other method. The total runtime in the table comprises ours and the baseline methods. Our approach boosts trees’ inlier ratio from 1.2% to 98%. All pipelines achieve 100% recall, with deviations limited to 0.27° and 0.05 m. TEASER and PCR-99 enhance the recall by 30% and 50%, respectively. TEASER, SC2PCR, MAC, and PCR-99 are 332 ms, 9 s, 330 s, and 10 s faster than their original methods, respectively. We boost the inlier ratio in sequence 01 jumps from 2.4% to 48%. TEASER, SC2PCR and MAC demonstrate 12%, 15%, and 10% recall improvement, respectively. TEASER, SC2PCR, MAC, and PCR-99 become faster by approximately 70 ms, 2 s, 30 s, and 12 s, respectively. The results prove that cascading our method with other methods effectively boosts registration performance and facilitates a qualitative leap in speed.
@@
IV-E Noise Study
As shown in Fig. 9, we evaluate the registration performance under various noise levels. In each, we randomly generate 3000 correspondences between -100 m and 100 m, satisfying the same transformation. Then, we replace 10%, 50%, 90%, 95%, and 98% correspondences with random values between -100 m and 100 m as outliers. Finally, we add 50 levels of Gaussian noise, with standard deviations $\sigma_{noise}$ from 0.1 m to 5 m, to the remaining inliers. We run 100 tests to compute recall, average rotation error, and translation error. Registration is successful if the RMSE of true inliers is less than three times $\sigma_{noise}$. The results show an overall recall of over 98%. With 98% outliers, the translation error stays below the noise level, and the rotation error remains under 2°, highlighting the method’s robustness.
(a) Recall
(b) Rotation and translation error
Figure 9: Registration across various noise levels. (b): Rotation (left) and translation (right) errors are shown in one figure.
IV-F Ablation Study
@@
TABLE V: Outlier removal analysis for three RANSAC modules. 1R, 2R, and 3R denote one-point, two-point, and three-point RANSAC, respectively. IR is the inlier ratio. IN means the inlier number.
@@ V Conclusion
We propose a two-stage consensus filtering method for correspondence-based registration. Our approach restores the RANSAC family to SOTA performance. By combining length consistency with angle consistency, we eliminate the majority of outliers, thus reducing the requisite iterations by three-point RANSAC. We combine three RASNAC modules and iterative weighted least squares into a complete registration pipeline that accelerates registration without compromising accuracy. Experimental results demonstrate a three-orders-of-magnitude speed enhancement while maintaining high registration quality. Moreover, integrating our method with existing approaches can notably improve efficiency and recall.